OpenStudy (anonymous):

log(logx)=2 Solve for x

OpenStudy (mathstudent55):

Use definition of log. Using base 10, log(x) = y means 10^y = x log(logx) = 2 means 10^2 = log x Use the definition again: log x = 10^2 means 10^(10^2) = x x = 10^(10^2) If the base is e instead of x, then pelace the 10's in the answer with e's

OpenStudy (anonymous):

log(logx)=2 Let both sides of the equation be the exponent of base 10. 10^(log(logx))=10^(2) When the base of a logarithm in the exponent and the base of the exponent are the same, the result is the argument of the logarithm. logx=10^(2) Squaring a number is the same as multiplying the number by itself (10*10). logx=10*10 Multiply 10 by 10 to get 100. logx=100 Let both sides of the equation be the exponent of base 10. 10^(logx)=10^(100) When the base of a logarithm in the exponent and the base of the exponent are the same, the result is the argument of the logarithm. x=10^(100) Raising a number to the 100th power is the same as multiplying the number by itself 100 times. x=1E+100

OpenStudy (mathstudent55):

*replace (not pelace)^

OpenStudy (anonymous):

Thank you all for the help. This looks easy now ;-)